Problem 10-04A (Video) At the beginning of 2018, Sheffield Company acquired equipment costing $175,600. It was estimated that this equipment would have a useful life of 6 years and a salvage value of $17,560 at that time. The straight-line method of depreciation was considered the most appropriate to use with this type of equipment. Depreciation is to be recorded at the end of each year.

During 2020 (the third year of the equipment’s life), the company’s engineers reconsidered their expectations, and estimated that the equipment’s useful life would probably be 7 years (in total) instead of 6 years. The estimated salvage value was not changed at that time. However, during 2023 the estimated salvage value was reduced to $5,000.

Indicate how much depreciation expense should be recorded each year for this equipment, by completing the following table.

Year Depreciation expense Accumulated expense

2018

2019

2020

2021

2022

2023

2024
Answer #1

Calculate depreciation expense

Year Depreciation expense Accumulated depreciation
2018 (175600-17560/6) = 26340 26340
2019 26340 52680
2020 (175600-52680-17560/5) = 21072 73752
2021 21072 94824
2022 21072 115896
2023 (175600-115896-5000/2) = 27352 143248
2024 27352 170600